Need help with what I got wrong, please.eBook B Print Que 6 Partially correct Mark 4.00 out of 10,00 P Flag question Edit question Compute, Disaggregate, and InterprNeed help with what I got wrong, please.
Answer #1

a) Return on assets= Net Income/ Average Total Assets

CVS Health Walgreen
Net Income(A) 6623 4101
Total Assets-2016(B) 94462 72688
Total Assets-2017(C) 95131 66009
Average Total Assets (D)=(B)+(C)/2 94796.5 69348.5
Return on assets=(A)/(D) 7% 5.9%

b) Profit Margin= [Net Profit+Interest Expenses(1-Tax rate)] /Sales Revenue

CVS Health Walgreen
Net Profit(A) 6623 4101
Interest(B) 1062 728
Interest*(1-Tax rate)(C) 690.3 473.2
Net profit + Intt(1-t)=(D)=(A)+(C) 7313.3 4574.2
Sales Revenue(E) 184765 118214
Profit Margin (D)/(E) 4% 3.9%

c) Return on Financial leverage= Return on Equity-Return on Assets

CVS Health Walgreen
ROE 17.8% 14%
ROA 7% 5.9%
ROFL 10.8% 8.1%

Please Rate if it helps you or let me know if anything else